# From What to Why: A Multi-Agent System for Evidence-based Chemical Reaction Condition Reasoning

### v1

@@ B.2 LLM-Based Multi-Agent Systems
To address hallucinations and unreliable reasoning, multi-agent collaboration has emerged as a promising direction. (Du et al., 2023) proposed a multi-agent debate framework where LLMs iteratively critique each other’s answers, leading to improved factuality and robustness. (Zhu et al., 2025) benchmarked agent interactions across collaborative and competitive settings, revealing that structured debate and agent role specialization improve task success. Recent work further explores coordination protocols. (Kaesberg et al., 2025) found that consensus-based decision-making outperforms majority voting on complex QA tasks, while (Zhang et al., 2024a) introduced a compression pipeline that reduces inter-agent communication by up to 70% without degrading performance. (Wu et al., 2025) introduced Agentic Reasoning, a general framework for LLMs to call sub-agents (e.g., web search, code execution, memory management), enabling long-horizon, tool-rich scientific workflows. Together, these systems demonstrate that combining LLMs with external tools, structured memory, and agent-level reasoning can produce scalable, verifiable pipelines for high-stakes domains. However, how to enhance the factuality and reliability of reaction condition prediction remains largely unexplored.
@@ C.1 Algorithm of ChemMAS Framework
This alternating scheme combines supervised teaching of tool protocols with reinforcement alignment for accuracy and collaboration, resulting in a robust tool-aware reasoning model $\pi^{\mathrm{RL}}_{\theta}$ with interpretable, consistent behavior.
@@
1:Initialize backbone model $\pi_{\theta}\leftarrow\text{Qwen3-8B-Instruct}$ $\triangleright$ AdamW ($\beta{=}(0.9,0.95)$), lr $2{\times}10^{-5}$, wd $0.1$, batch 128128
@@
Stage II: Tool Incentivization (RL with GRPO) /* align accuracy & tool use */
@@ D.1 Training Pipeline
The input is limited to Reactant and Product SMILES, and the output is the predicted Reaction Type. During SFT, the supervision target is structured as a step-wise chain that explicitly serializes three tool invocations—Functional Group Tagger, Constraint Engine, and Chemical Knowledge Base Searching—before emitting the final reaction type. This design enables the model to learn when and how to call tools. In the subsequent RL stage, we apply a hierarchical reward that integrates format correctness, answer accuracy, and collaborative multi-tool usage.
@@ D.2 Evaluation Details
We evaluate general-purpose LLMs in a controlled candidate-ranking regime aligned with the ChemMAS pipeline. Directly prompting models with only Reactant and Product SMILES yields an excessively large decision space, leading to chemically plausible yet inaccurate suggestions and a Top-1 accuracy of approximately 5%. To obtain a faithful assessment, for each reaction a high-recall pool is first constructed via Multi-Channel Recall—aggregating reaction-base retrieval, functional-group cues, constraint heuristics, and memory lookup—to produce a Top-5000 candidate set spanning Catalyst, Solvent1, Solvent2, Reagent1, and Reagent2. Each model ranks within the same 5k pool and outputs a Top-50 list per head. All models receive identical candidate sets, instructions, and judgment interfaces, and are not permitted to modify the pool, ensuring that differences reflect discriminative ranking and evidence integration rather than retrieval coverage. This protocol mitigates search-space inflation, reduces hallucination, and provides an evaluation setting consistent with the workflow of the framework.
D.3 Prompt Templates
@@
Appendix E Results and Discussions
E.1 Additional Quantitative Results
@@
As shown in Figure 7, introducing specialized agents consistently improves Top-5 Accuracy over the $\mathcal{A}_{Gen}+\mathcal{A}_{Full}$ baseline. $\mathcal{A}_{Cat}$ delivers targeted gains on Catalyst (+10.1%), aligning with its role specialization. $\mathcal{A}_{Sol}$ contributes the most on solvents, improving Solvent1 and Solvent2 by +16.4% and +13.4%, respectively. $\mathcal{A}_{Rea}$ yields the largest boosts on reagents (e.g., Reagent1/2 with gains around +18.7% and +13.9%). When specialized agents are combined (e.g., +Cat+Sol, +Sol+Rea, +Cat+Rea), the improvements remain additive and stable across condition types, and the Full System shows the most consistent Top-5 lift across all five heads, indicating effective collaboration among role-specialized experts.
@@
As shown in Figure 8, the same trend holds for Top-10 Accuracy. $\mathcal{A}_{Cat}$ most strongly benefits Catalyst (+13.1%). $\mathcal{A}_{Sol}$ provides clear gains on Solvent1/2 (e.g., +10.8% and +13.6%). $\mathcal{A}_{Rea}$ again dominates on Reagent1/2 with sizeable increments (e.g., +17.2% and +9.8%). Pairwise combinations further enhance coverage across heads, and the Full System achieves the highest Top-10 metrics in a macro sense, evidencing that multi-agent collaboration scales beyond single-head expertise and produces robust gains under larger candidate sets.
Figure 7: Multi-agent ablation: Top-5 accuracy improvements across Catalyst, Solvent1/2, and Reagent1/2 when adding specialized agents on top of $\mathcal{A}_{Gen}$$\mathcal{A}_{Full}$.
Figure 8: Multi-agent ablation: Top-10 accuracy improvements across Catalyst, Solvent1/2, and Reagent1/2 when adding specialized agents on top of $\mathcal{A}_{Gen}$$\mathcal{A}_{Full}$.
E.2 Result Visualization
@@
Overall, the visualization confirms that the framework not only achieves high top-$k$ accuracy but also produces predictions that are chemically interpretable and robust. The ability to provide both exact matches and plausible alternatives underscores the potential of our approach for assisting chemists in condition selection and experimental design.

### v2

@@ B.2 LLM-Based Multi-Agent Systems
To address hallucinations and unreliable reasoning, multi-agent collaboration has emerged as a promising direction. (Du et al., 2023) proposed a multi-agent debate framework where LLMs iteratively critique each other’s answers, leading to improved factuality and robustness. (Zhu et al., 2025) benchmarked agent interactions across collaborative and competitive settings, revealing that structured debate and agent role specialization improve task success. Recent work further explores coordination protocols. (Kaesberg et al., 2025; Yang et al., 2026) found that consensus-based decision-making outperforms majority voting on complex QA tasks, while (Zhang et al., 2024a) introduced a compression pipeline that reduces inter-agent communication by up to 70% without degrading performance. (Wu et al., 2025) introduced Agentic Reasoning, a general framework for LLMs to call sub-agents (e.g., web search, code execution, memory management), enabling long-horizon, tool-rich scientific workflows. Together, these systems demonstrate that combining LLMs with external tools, structured memory, and agent-level reasoning can produce scalable, verifiable pipelines for high-stakes domains. However, how to enhance the factuality and reliability of reaction condition prediction remains largely unexplored.
@@ C.1 Algorithm of ChemMAS Framework
This alternating scheme combines supervised teaching of tool protocols with reinforcement alignment for similarity and collaboration, resulting in a robust tool-aware reasoning model $\pi^{\mathrm{RL}}_{\theta}$ with interpretable, consistent behavior.
@@
1:Initialize backbone model $\pi_{\theta}\leftarrow\text{Qwen3-8B-Instruct}$ $\triangleright$ AdamW ($\beta{=}(0.9,0.95)$), lr $2{\times}10^{-5}$, wd $0.1$, batch $128$
@@
Stage II: Tool Incentivization (RL with GRPO) /* align similarity & tool use */
@@ D.1 Training Pipeline
The input is limited to Reactant and Product SMILES, and the output is the predicted Reaction Type. During SFT, the supervision target is structured as a step-wise chain that explicitly serializes three tool invocations—Functional Group Tagger, Constraint Engine, and Chemical Knowledge Base Searching—before emitting the final reaction type. This design enables the model to learn when and how to call tools. In the subsequent RL stage, we apply a hierarchical reward that integrates format correctness, answer similarity, and collaborative multi-tool usage.
@@ D.2 Evaluation Setting Details
We evaluate general-purpose LLMs in a controlled candidate-ranking regime aligned with the ChemMAS pipeline. Directly prompting models with only Reactant and Product SMILES yields an excessively large decision space, leading to chemically plausible yet inaccurate suggestions and a Top-1 similarity of approximately 5%. To obtain a faithful assessment, for each reaction a high-recall pool is first constructed via Multi-Channel Recall—aggregating reaction-base retrieval, functional-group cues, constraint heuristics, and memory lookup—to produce a Top-5000 candidate set spanning Catalyst, Solvent1, Solvent2, Reagent1, and Reagent2. Each model ranks within the same 5k pool and outputs a Top-50 list per head. All models receive identical candidate sets, instructions, and judgment interfaces, and are not permitted to modify the pool, ensuring that differences reflect discriminative ranking and evidence integration rather than retrieval coverage. This protocol mitigates search-space inflation, reduces hallucination, and provides an evaluation setting consistent with the workflow of the framework.
Appendix E Evaluation Protocol Details
In this section, we provide the formal definition of the structure-aware evaluation metrics used in our experiments. The Reaction Condition Recommendation (RCR) task requires models to predict appropriate reaction conditions given reactants and products. We evaluate the quality of predicted condition SMILES strings against ground truth annotations using molecular fingerprint similarity metrics.
E.1 SMILES Canonicalization and Validity
Prior to fingerprint calculation, all SMILES strings undergo canonicalization to ensure consistent molecular representations. Let $\hat{s}$ be the predicted SMILES and $s^{*}$ be the ground truth SMILES. The canonicalization procedure converts input SMILES to a standardized canonical form:
$s_{\text{canonical}}=\text{Canonicalize}(s_{\text{input}})$ (13)
This process removes representational ambiguity. Consistent with our evaluation constraints, stereochemical information is excluded during canonicalization (isomericSmiles=False) to focus evaluation on constitutional structure.
The validity metric quantifies the proportion of predictions that correspond to chemically valid molecular structures:
$\text{Validity}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}\left[\text{IsValid}(\hat{s}_{i})\right]$ (14)
where $\mathbb{1}[\cdot]$ is the indicator function and $\text{IsValid}(\cdot)$ returns true if the SMILES string can be successfully parsed into a valid molecular graph by RDKit (version 2023.03 or later). Invalid predictions are assigned a similarity score of 0 for all fingerprint metrics.
E.2 Tanimoto Similarity
All fingerprint-based similarity calculations employ the Tanimoto coefficient (Jaccard index). For two molecular fingerprints represented as bit vectors $\mathbf{A}$ and $\mathbf{B}$ (corresponding to the predicted molecule $M_{p}$ and ground truth molecule $M_{g}$), the Tanimoto similarity is defined as:
$T(\mathbf{A},\mathbf{B})=\frac{|\mathbf{A}\cap\mathbf{B}|}{|\mathbf{A}\cup\mathbf{B}|}=\frac{c}{a+b-c}$ (15)
where $a$ denotes the number of bits set to 1 in $\mathbf{A}$, $b$ denotes the number of bits set to 1 in $\mathbf{B}$, and $c$ denotes the number of bits set to 1 in both fingerprints simultaneously.
E.3 Molecular Fingerprint Types
We employ three complementary molecular fingerprint representations to capture different aspects of molecular structure:
RDK Topological Fingerprint ($f_{\text{RDK}}$).
This is a path-based topological fingerprint. The algorithm enumerates all linear paths of length $l\in[1,7]$ atoms within the molecular graph. Each path is encoded as a hash incorporating atomic numbers, bond types, and connectivity. The resulting hash values are mapped to a bit vector of length 2048. The similarity is computed as:
$S_{\text{RDK}}(M_{p},M_{g})=T\left(f_{\text{RDK}}(M_{p}),f_{\text{RDK}}(M_{g})\right)$ (16)
MACCS Keys Fingerprint ($f_{\text{MACCS}}$).
The MACCS keys fingerprint consists of 166 predefined structural keys, each corresponding to specific substructures (e.g., hydroxyl, carbonyl, aromatic rings). For each key $k_{i}$ ($i\in[1,166]$), the bit is set to 1 if the substructure is present. This metric is particularly valuable for comparing molecules based on functional group composition:
$S_{\text{MACCS}}(M_{p},M_{g})=T\left(f_{\text{MACCS}}(M_{p}),f_{\text{MACCS}}(M_{g})\right)$ (17)
Morgan Circular Fingerprint ($f_{\text{Morgan}}$).
We employ Morgan fingerprints with a radius parameter of $r=2$, equivalent to ECFP4. This algorithm captures the local chemical environment by iteratively identifying atom identifiers and their neighbors. It excels at detecting localized structural differences and functional group modifications:
$S_{\text{Morgan}}(M_{p},M_{g})=T\left(f_{\text{Morgan}}(M_{p},r{=}2),f_{\text{Morgan}}(M_{g},r{=}2)\right)$ (18)
E.4 Aggregate Evaluation Metrics
To assess the model performance, we utilize the Fingerprint Tanimoto Score (FTS), defined as the validity-weighted average of the three fingerprint similarities:
$\text{FTS}=\left(\frac{\bar{S}_{\text{RDK}}+\bar{S}_{\text{MACCS}}+\bar{S}_{\text{Morgan}}}{3}\right)\times\text{Validity}$ (19)
where $\bar{S}_{\text{type}}$ represents the mean similarity score across all test samples for that specific fingerprint type.
E.5 Top-$k$ Similarity
Since our model generates a ranked list of candidate predictions, we also report a Top-$k$ metric. Let $\mathcal{S}(\hat{s},s^{*})$ be the pairwise similarity for a single instance, defined as the average of the three fingerprint Tanimoto coefficients (assigned as 0 if $\hat{s}$ is invalid).
For a dataset of $N$ reaction instances, where $s^{*}_{i}$ is the ground-truth SMILES and $\hat{S}_{i,k}=\{\hat{s}_{i,1},\dots,\hat{s}_{i,k}\}$ is the set of the top-$k$ predicted SMILES strings:
$\text{Score}@k=\frac{1}{N}\sum_{i=1}^{N}\max_{j\in\{1,\dots,k\}}\mathcal{S}(\hat{s}_{i,j},s^{*}_{i}).$ (20)
Figure 6: Heavy-atom and molecular-weight distributions for the Train, Validation, and Test sets (top to bottom). Left column: heavy atom count; right column: molecular weight (g/mol).
Appendix F Dataset Details
Public Dataset.
We use the RCR subset of ChemCoTBench, which contains 90 well-structured samples covering 10 reaction types. For each reaction type, there are 9 examples: 3 focused on catalyst prediction, 3 on reagent prediction, and 3 on solvent prediction. All chemical entities (reactants/products/conditions) are represented in SMILES format to ensure consistency.
Private Dataset.
We curate a large-scale private dataset of organic reactions to supplement existing public benchmarks and better represent real-world experimental scenarios. Sourced from the internal database of an anonymous chemical research institution, this dataset is rigorously digitized and structured, comprising 544,591 high-quality entries. Similar in nature to the USPTO-condition dataset, it encompasses a broad spectrum of known chemical reactions, reflecting a chemical space.
For data standardization, all chemical entities are represented in SMILES format. Each SMILES string is processed with RDKit to construct a molecular graph; unparseable strings are discarded. For every valid molecule, we compute the total atom count, the heavy-atom count (all non-hydrogen atoms), the molecular weight as the sum of average atomic masses (g/mol), and the exact mass as the sum of isotopic masses (g/mol). We then analyze and visualize the heavy-atom and molecular-weight distributions for the training, validation, and test sets, where the left column shows the heavy-atom counts and the right column shows the molecular weights in g/mol, as shown in Figure 6. We frame the task as a reaction condition prediction problem: for each entry, the reactants and products serve as the input, while the reaction conditions are defined as the output. To enable fine-grained prediction, the output is structured into five distinct roles: catalyst (Catalyst1), solvents (Solvent1, Solvent2), and reagents (Reagent1, Reagent2).
Based on this setting, we construct Question-Answer (QA) pairs to facilitate model training. The dataset is randomly split into training, validation, and test sets with a ratio of 8:1:1. The inclusion of this private dataset provides robust supervision signals and allows for the evaluation of model generalization in complex, realistic chemical contexts.
Appendix G Prompt Templates
@@
Appendix H Results and Discussions
H.1 Additional Quantitative Results
@@
As shown in Figure 9, introducing specialized agents consistently improves Top-5 Similarity over the $\mathcal{A}_{Gen}+\mathcal{A}_{Full}$ baseline. $\mathcal{A}_{Cat}$ delivers targeted gains on Catalyst (+10.1%), aligning with its role specialization. $\mathcal{A}_{Sol}$ contributes the most on solvents, improving Solvent1 and Solvent2 by +16.4% and +13.4%, respectively. $\mathcal{A}_{Rea}$ yields the largest boosts on reagents (e.g., Reagent1/2 with gains around +18.7% and +13.9%). When specialized agents are combined (e.g., +Cat+Sol, +Sol+Rea, +Cat+Rea), the improvements remain additive and stable across condition types, and the Full System shows the most consistent Top-5 lift across all five heads, indicating effective collaboration among role-specialized experts.
@@
As shown in Figure 10, the same trend holds for Top-10 Similarity. $\mathcal{A}_{Cat}$ most strongly benefits Catalyst (+13.1%). $\mathcal{A}_{Sol}$ provides clear gains on Solvent1/2 (e.g., +10.8% and +13.6%). $\mathcal{A}_{Rea}$ again dominates on Reagent1/2 with sizeable increments (e.g., +17.2% and +9.8%). Pairwise combinations further enhance coverage across heads, and the Full System achieves the highest Top-10 metrics in a macro sense, evidencing that multi-agent collaboration scales beyond single-head expertise and produces robust gains under larger candidate sets.
Figure 9: Multi-agent ablation: Top-5 similarity improvements across Catalyst, Solvent1/2, and Reagent1/2 when adding specialized agents on top of $\mathcal{A}_{Gen}$$\mathcal{A}_{Full}$.
Figure 10: Multi-agent ablation: Top-10 similarity improvements across Catalyst, Solvent1/2, and Reagent1/2 when adding specialized agents on top of $\mathcal{A}_{Gen}$$\mathcal{A}_{Full}$.
H.2 Result Visualization
@@
Overall, the visualization confirms that the framework not only achieves high top-$k$ similarity but also produces predictions that are chemically interpretable and robust. The ability to provide both exact matches and plausible alternatives underscores the potential of our approach for assisting chemists in condition selection and experimental design.